OpenStudy (anonymous):

HELP find x: 1+3(cos(x))^2=4sin(pi/2+x)

OpenStudy (loser66):

manipulate the RHS 4sin(pi/2 +x) = 4cosx replace to get \[1+3cos^2x=4cosx\] \[3cos^2x-4cosx+1=0\] solve as quadratic equation with variable is cos x after getting values of cos x plug back trig to solve for x